Ops/Comm Scheduling meeting¶
Location : CR23 in R1 @ 15:30 pm on August 23rd
Phone access : ReadyTalk number
Access Code: 2155402
- DC report
- Plans for the following week
- http://micewww.pp.rl.ac.uk/projects/operations/wiki/_StepIV_Shutdown_Summer_ (We've gone through this at the MEMO today. I don't propose to go through it again)
- DONM : 30th August in CR21 in R1
Present : SB, AG, MC, VB, HN, PH, CW, JC
- AC work and finishing the water work dominated last week's work.
- Another leak in one of the AC units has been found and will be fixed tomorrow. Sometime tomorrow all the AC work will be done and we can lock up.
- Tracker cryos 3/4 at Hall temperature due to the water switchover. Alan wants some checks made before the cryos are cooled down again.
- Water changeover on the roof is taking place now. Earlier today a valve was closed which shouldn't have been closed. The water temp rose to 90 degrees!
and water started filling the trench. We've taken several hours to recover from this. Magnets were valved off so magnets are OK.
- Vicky attempted to bring up the DS Linde compressor but has found a leak. The low pressure circuit won't hold pressure. Mike suspects a leak. The servicing
company will come tomorrow and fix it.
- Action Are ISIS meant to touch the compressors? Do we need an isolation procedure for ISIS to follow if they are trying to cut the water? Talk to Colin/John about shared services?
- Henry has inspected the target in the synch and everything looks OK. Target is ready to go.
- Action Josef has done 51% of the spectrometer documentation.
Plans for the following week¶
- Wednesday : Finish AC work and if the water is OK, closed the Hall up ready for the magnet tests.
- Thursday/Friday : power SSU and look at movement.
- Friday : power all coils to 3T.
- Action Paul will send a plan for the next week.
- Action Get some form of neutron monitor in DSA (Henry)